# Revisiting Synthetic Human Trajectories: Imitative Generation and Benchmarks Beyond Datasaurus

### v1

@@ Abstract.
Human trajectory data, which plays a crucial role in various applications such as crowd management and epidemic prevention, is challenging to obtain due to practical constraints and privacy concerns. In this context, synthetic human trajectory data is generated to simulate as close as possible to real-world human trajectories, often under summary statistics and distributional similarities. However, the complexity of human mobility patterns is oversimplified by these similarities (a.k.a. “Datasaurus”), resulting in intrinsic biases in both generative model design and benchmarks of the generated trajectories. Against this background, we propose MIRAGE, a huMan-Imitative tRAjectory GenErative model designed as a neural Temporal Point Process integrating an Exploration and Preferential Return model. It imitates the human decision-making process in trajectory generation, rather than fitting any specific statistical distributions as traditional methods do, thus avoiding the Datasaurus issue. Moreover, we also propose a comprehensive task-based evaluation protocol beyond Datasaurus to systematically benchmark trajectory generative models on four typical downstream tasks, integrating multiple techniques and evaluation metrics for each task, to comprehensively assess the ultimate utility of the generated trajectories. We conduct a thorough evaluation of MIRAGE on three real-world user trajectory datasets against a sizeable collection of baselines. Results show that compared to the best baselines, MIRAGE-generated trajectory data not only achieves the best statistical and distributional similarities with 59.0-71.5% improvement, but also yields the best performance in the task-based evaluation with 10.9-33.4% improvement.
Human Mobility, Trajectory, Generative Model, Mobility Simulation
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Spatial-temporal systems
@@ 1. Introduction
Human trajectory data is a key ingredient for a wide range of applications, including urban planning (Zhou et al., 2018), traffic management (Naboulsi et al., 2016), epidemic analysis (Gao et al., 2020), predictive policing (Wu et al., 2022), and location-based services (Gao and Liu, 2015). These applications heavily rely on the quality of human mobility models learnt from human trajectory data. However, acquiring large-scale real human trajectory data is often challenging due to practical constraints and privacy concerns (Kim et al., 2020). Therefore, synthetic human trajectory data has been widely used as an alternative, where generative models are learnt to generate artificial trajectories that closely resemble real-world human trajectories, making human trajectory data more readily available.
In the current literature, existing works on synthetic human trajectory generation mostly focus on the resemblance under summary statistics and distributional similarities (Kapp et al., 2023). For example, these works measure the resemblance in different aspects of mobility characteristics such as spatial distribution (e.g. G-Rank) (Jiang et al., 2023), temporal distribution (e.g. stay duration) (Pappalardo and Simini, 2018), OD flows (trips per OD pair) (Yuan et al., 2011), and user mobility patterns (e.g. I-Rank and DailyLoc) (Xu et al., 2021), using divergence/distance metrics such as Kullback-Leibler divergence (KLD) (Bindschaedler and Shokri, 2016), Jensen-Shannon divergence (JSD) (Ouyang et al., 2018; Feng et al., 2020), earth mover’s distance (EMD) (Anda et al., 2021), Root Mean Squared Error (RMSE) (Pappalardo and Simini, 2018; Badu-Marfo et al., 2022). These similarities often serve on one hand as part of the model design such as the model fitting objective (Jiang et al., 2016), while on the other hand also as the benchmarks for evaluating trajectory generative models (Long et al., 2023; Ouyang et al., 2018; Huang et al., 2019; Xiong et al., 2023). However, while these similarities provide insights into the differences between real and generated data from various perspectives, they indeed oversimplify the complexity of human mobility patterns, resulting in intrinsic biases in both generative model design and benchmarks of the generated trajectories. Specifically, datasets that are similar over a number of statistical properties may yield very different patterns, known as “Datasaurus” (Cairo, 2016; Matejka and Fitzmaurice, 2017). In the context of trajectory data, it implies that statistically/distributionally similar trajectories may imply different mobility patterns, which thus leads to different performances in downstream tasks (as evidenced by our experiments in Section 5.3). Therefore, the resemblance under summary statistics and distributional similarities cannot fully reflect the ultimate utility of generated trajectories in supporting downstream tasks.
@@
To address the above issues, we propose MIRAGE, a huMan-Imitative tRAjectory GenErative model, which imitates the holistic human decision-making process in trajectory generation and does not explicitly fit any specific statistical distributions (while traditional EPR model does), thus avoiding the “Datasaurus” issue. Specifically, a human trajectory consisting of a sequence of stochastic presence events on continuous time is naturally a Temporal Point Process (TPP). Subsequently, we design MIRAGE as a neural TPP with intensity-free parameterization, benefiting from both the flexibility of the neural network encoding the trajectory history and the efficiency of the TPP in modeling continuous-time stochastic events (Shchur et al., 2021b). To generate one event of a human trajectory, we first sample an event time from the neural TPP based on the encoded trajectory history, and then sample the activity category further conditioned on the sampled event time. Afterward, we design a neural EPR model, mimicking the human decision-making process to choose either exploring unvisited locations or returning to previously visited locations, conditioned on the sampled time and activity category. We also adopt a user variational autoencoder to imitate individual preferences in the trajectory generation process.
@@
  • We conduct a thorough evaluation of MIRAGE on three real-world human trajectory datasets against a sizeable collection of state-of-the-art baselines. Results show that compared to the best baselines, MIRAGE-generated trajectory data not only achieves the best statistical and distributional similarities with 59.0-71.5% improvement but also yields the best performance in the task-based evaluation with 10.9-33.4% improvement.
@@ 2.1. Mobility Trajectory Generation
Early works mostly model individual mobility with explicit physical meanings (Gonzalez et al., 2008; Song et al., 2010b) and generate synthetic trajectories under the distributions of key characteristics observed in real mobility patterns, such as trip lengths, start locations, or start times, etc. (Roy et al., 2016; Mir et al., 2013). As a widely recognized mobility model, the Exploration and Preferential Return model (EPR) (Song et al., 2010a) unifies exploration and return mobility patterns by selecting new locations based on a random walk process for exploration and revisiting previously visited locations based on their frequency for preferential return. Subsequent studies further extend the EPR model by integrating sophisticated spatial or social information, such as mining the correlation between location capacity and social network sizes (Alessandretti et al., 2018), incorporating a nested gravity model into the EPR model (Pappalardo et al., 2015), linking mobility to social ties and studying the dynamics of spatial choices based on social behavior (Toole et al., 2015), and integrating the circadian propensity of human mobility and Markov-based models into the EPR model (Jiang et al., 2016). In addition, the EPR model is also shown to be universal in human behavior modeling in general, such as user activities in recommendation systems (Sun et al., 2020; Onuma et al., 2009) and human behaviors in cyberspace (Hu et al., 2019; Zhao et al., 2014; Hu et al., 2018). However, these models, often reliant on heuristic statistical assumptions, have limitations in accurately modeling the complex mobility patterns observed in real-world trajectories.
@@ 2.2. Synthetic Mobility Trajectory Benchmarks
The benchmarks for synthetic mobility trajectories can be classified into two categories (Kapp et al., 2023). First, statistical and distributional similarities are the most widely used benchmarks, such as Kullback-Leibler divergence (KLD) (Bindschaedler and Shokri, 2016), Jensen-Shannon divergence (JSD) (Ouyang et al., 2018; Feng et al., 2020), earth mover’s distance (EMD) (Anda et al., 2021), Root Mean Squared Error (RMSE) (Pappalardo and Simini, 2018; Badu-Marfo et al., 2022), which are used to measure the similarities between real and generated trajectories in different aspects. For example, typical mobility statistics include the radius of gyration (Feng et al., 2020; Jiang et al., 2023; Yuan et al., 2022), the number of distinct locations visited per user per day (Feng et al., 2020; Xu et al., 2021), I-Rank (frequency of visiting personal top locations) (Feng et al., 2020; Pappalardo and Simini, 2018), and the number of daily trips per user (Pappalardo and Simini, 2018), trip lengths between consecutive trajectory points or between the origin and destination (Feng et al., 2020; Gursoy et al., 2018; Jiang et al., 2023; Sun et al., 2023; Xiong et al., 2023); spatial distributions characterize the distribution of locations based on factors like visits per location (i.e. G-Rank) (Feng et al., 2020; Jiang et al., 2023; Ouyang et al., 2018; Wang et al., 2021; Xiong et al., 2023) or location popularity ranking (Du et al., 2023; Gursoy et al., 2018); temporal distributions characterize the number of trips per hour of the day (Pappalardo and Simini, 2018), stay duration (Feng et al., 2020; Long et al., 2023; Ouyang et al., 2018) and time intervals between check-ins (Yuan et al., 2022). However, while these similarity metrics provide insights into the differences between real and generated data from various perspectives, they cannot fully reflect the ultimate utility of generated trajectories in supporting downstream tasks.
Second, benchmarking on downstream tasks recently emerged as an evaluation scheme for synthetic mobility trajectories. These tasks include road map updating (Cao and Li, 2021), next location prediction (Jiang et al., 2023; Wei et al., 2021; Long et al., 2023), and spreading simulation (Xiong et al., 2023; Feng et al., 2020; Yuan et al., 2022). However, these works often use a heuristically designed downstream task with one specific technique/algorithm to solve the task, which leads to unknown biases in the utility evaluation. As evidenced in our experiments in Appendix A, the performance of different techniques solving the same task varies; heuristically choosing the results of one technique as the benchmark is thus untrustworthy. Therefore, we propose a comprehensive task-based evaluation protocol to systematically benchmark synthetic mobility trajectories.
@@ 3.2.2. TPP Parameterization
Traditional TPPs often specify a simple parametric intensity function capturing relatively simple patterns in event occurrences, such as the Hawkes process (Hawkes, 1971), which often leads to poor results because of its limited flexibility in modeling complex data. In this context, neural TPPs are developed to use neural networks to learn complex dependencies from the history of TPPs, and then approximate the intensity function by some parametric forms (Shchur et al., 2021a). However, these intensity-based approaches cannot achieve flexibility, efficiency, and ease-to-use simultaneously (Shchur et al., 2019). Alternatively, an intensity-free learning model of neural TPPs is proposed to use a mixture of log-normal distributions to represent the conditional probability density function $p^{*}(\tau)$ directly, and thus bypass the cumbersome intensity function (Shchur et al., 2019). Due to its advantage of closed-form sampling and likelihood computation, the intensity-free approach has been widely adopted in learning neural TPPs (Xu et al., 2021; Liu et al., 2023; Gupta et al., 2022; Feng et al., 2023). We also adopt the intensity-free approach in this work.
@@ 4.3.3. Location decoder with neural EPR
Location sampling in the return mode. In the return mode $z=return$, we consider the temporal periodicity of human behaviors, where individuals tend to revisit previously visited locations under a regular temporal distance, such as returning home daily (with a temporal distance of 24 hours) (Song et al., 2010a). Therefore, the return mode is designed to imitate such revisiting behaviors. Specifically, instead of imposing specific constraints or predefined temporal periodicity as (Yang et al., 2020a), we use temporal distance embeddings to learn the probability distribution of returning to previously visited locations. We define the temporal distance under an hour granularity between the $(i+1)$-th and $j$-th events as $\Delta^{t}_{i+1,j}=|t_{i+1}-t_{j}|\in\mathbb{Z}_{\geq 0}$. Subsequently, for the next event time $t_{i+1}$, we obtain its temporal distances to all previous events as follows:
@@
Note that here $p^{*}(l_{i+1}|\tau_{i+1},return)$ is not conditioned on the sampled next event category $k_{i+1}$ because the revisiting behavior depends only on the next event time. We then sample a previously visited location $l$ according to $p^{*}(l_{i+1}|\tau_{i+1},return)$. Our experiments also reveal that this learnt probability indeed exhibits a clear daily return pattern over different time distances (see Appendix B).
@@ 4.4.1. Model Training
In the training process, the above conditional probability distribution is conditioned on the real data without sampling. After training this objective function, we then train the user VAE on the learnt individual preference embeddings. The complexity of our model is discussed in Appendix C.
@@ 5.1.2. Baselines
We consider the following state-of-the-art baselines of three categories: statistical models Semi-Markov (Korolyuk et al., 1975) and TimeGeo (Jiang et al., 2016); neural TPP RMTPP (Du et al., 2016), ERTPP (Xiao et al., 2017), THP (Zuo et al., 2020), and LogNormMix (Shchur et al., 2019); deep learning generative models LSTM (Huang et al., 2015), SeqGAN (Yu et al., 2017), MoveSim (Feng et al., 2020), VOLUNTEER (Long et al., 2023) and DiffTraj (Zhu et al., 2024). The details of these baselines are in Appendix D.
@@ 5.1.3. Statistical and Distributional Similarity Metrics
Following previous works (Feng et al., 2020; Ouyang et al., 2018), we adopt popular metrics to evaluate the resemblance between real and generated trajectories in different aspects. Distance measures the distance between successive locations in a trajectory. Radius of gyration is calculated as the root mean squared distance of all locations from the central one in a trajectory. Interval is computed as time intervals between successive events in a trajectory. DailyLoc computes the unique locations visited by users. Category computes the overall distribution of the POI categories. We utilize the Jensen-Shannon divergence (JSD) (Fuglede and Topsoe, 2004) as the similarity metric between the distributions of real and generated trajectories.

### v2

@@ Abstract.
Human trajectory data, which plays a crucial role in various applications such as crowd management and epidemic prevention, is challenging to obtain due to practical constraints and privacy concerns. In this context, synthetic human trajectory data is generated to simulate as close as possible to real-world human trajectories, often under summary statistics and distributional similarities. However, these similarities oversimplify complex human mobility patterns (a.k.a. “Datasaurus”), resulting in intrinsic biases in both generative model design and benchmarks of the generated trajectories. Against this background, we propose MIRAGE, a huMan-Imitative tRAjectory GenErative model designed as a neural Temporal Point Process integrating an Exploration and Preferential Return model. It imitates the human decision-making process in trajectory generation, rather than fitting any specific statistical distributions as traditional methods do, thus avoiding the Datasaurus issue. We also propose a comprehensive task-based evaluation protocol beyond Datasaurus to systematically benchmark trajectory generative models on four typical downstream tasks, integrating multiple techniques and evaluation metrics for each task, to assess the ultimate utility of the generated trajectories. We conduct a thorough evaluation of MIRAGE on three real-world user trajectory datasets against a sizeable collection of baselines. Results show that compared to the best baselines, MIRAGE-generated trajectory data not only achieves the best statistical and distributional similarities with 59.0-67.7% improvement, but also yields the best performance in the task-based evaluation with 10.9-33.4% improvement. A series of ablation studies also validate the key design choices of MIRAGE. Our code and datasets are at https://github.com/UM-Data-Intelligence-Lab/MIRAGE.
Human Mobility; Trajectory; Generative Model; Simulation
ccs: Information systems Spatial-temporal systems
@@ 1. Introduction
Human trajectory data is a key ingredient for a wide range of applications, including urban planning (Zhou et al., 2018), traffic management (Naboulsi et al., 2016), epidemic analysis (Gao et al., 2020), predictive policing (Yang et al., 2018), and crowd monitoring (Chen et al., 2024). These applications heavily rely on the quality of human mobility models learnt from human trajectory data. However, acquiring large-scale real human trajectory data is often challenging due to practical constraints and privacy concerns. Therefore, synthetic human trajectory data has been widely used as an alternative, where generative models are learnt to generate artificial trajectories that closely resemble real-world human trajectories, making human trajectory data more readily available.
In the current literature, existing works on synthetic human trajectory generation mostly focus on the resemblance under summary statistics and distributional similarities (Kapp et al., 2023). For example, these works measure the resemblance in different aspects of mobility characteristics such as spatial distribution (e.g. G-Rank) (Jiang et al., 2023), temporal distribution (e.g. stay duration) (Pappalardo and Simini, 2018), OD flows (trips per OD pair) (Yuan et al., 2011), and user mobility patterns (e.g. I-Rank and DailyLoc) (Xu et al., 2021), using divergence/distance metrics such as Kullback-Leibler divergence (KLD) (Bindschaedler and Shokri, 2016), Jensen-Shannon divergence (JSD) (Ouyang et al., 2018; Feng et al., 2020), earth mover’s distance (EMD) (Anda et al., 2021), Root Mean Squared Error (RMSE) (Pappalardo and Simini, 2018). These similarities often serve on one hand as part of the model design such as the model fitting objective (Jiang et al., 2016), while on the other hand also as the benchmarks for evaluating trajectory generative models (Long et al., 2023; Ouyang et al., 2018; Huang et al., 2019; Xiong et al., 2023). However, while these similarities provide insights into the differences between real and generated data from various perspectives, they indeed oversimplify the complexity of human mobility patterns, resulting in intrinsic biases in both generative model design and benchmarks of the generated trajectories. Specifically, datasets that are similar over a number of statistical properties may yield very different patterns, known as “Datasaurus” (Cairo, 2016; Matejka and Fitzmaurice, 2017). In the context of trajectory data, it implies that statistically/distributionally similar trajectories may imply different mobility patterns, which thus leads to different performances in downstream tasks (as evidenced by our experiments in Section 5.3). Therefore, the resemblance under summary statistics and distributional similarities cannot fully reflect the ultimate utility of generated trajectories in supporting downstream tasks.
@@
To address the above issues, we propose MIRAGE, a huMan-Imitative tRAjectory GenErative model, which imitates the holistic human decision-making process in trajectory generation and does not explicitly fit any specific statistical distributions (while traditional EPR model does), thus avoiding the “Datasaurus” issue. Specifically, a human trajectory consisting of a sequence of stochastic presence events on continuous time is naturally a Temporal Point Process (TPP). Subsequently, we design MIRAGE as a neural TPP with intensity-free parameterization, benefiting from both the flexibility of the neural network encoding the trajectory history and the efficiency of the TPP in modeling continuous-time stochastic events (Shchur et al., 2021). To generate one event of a human trajectory, we first sample an event time from the neural TPP based on the encoded trajectory history, and then sample the activity category further conditioned on the sampled event time. Afterward, we design a neural EPR model, mimicking the human decision-making process to choose either exploring unvisited locations or returning to previously visited locations, conditioned on the sampled time and activity category. We also adopt a user variational autoencoder to imitate individual preferences in the trajectory generation process.
@@
  • We conduct a thorough evaluation of MIRAGE on three real-world human trajectory datasets against a sizeable collection of state-of-the-art baselines. Results show that compared to the best baselines, MIRAGE-generated trajectory data not only achieves the best statistical and distributional similarities with 59.0-67.7% improvement but also yields the best performance in the task-based evaluation with 10.9-33.4% improvement.
@@ 2.1. Mobility Trajectory Generation
Early works mostly model individual mobility with explicit physical meanings (Gonzalez et al., 2008; Song et al., 2010b) and generate synthetic trajectories under the distributions of key characteristics observed in real mobility patterns, such as trip lengths, start locations, or start times, etc. As a widely recognized mobility model, the Exploration and Preferential Return model (EPR) (Song et al., 2010a) unifies exploration and return mobility patterns by selecting new locations based on a random walk process for exploration and revisiting previously visited locations based on their frequency for preferential return. Subsequent studies further extend the EPR model by integrating sophisticated spatial or social information, such as mining the correlation between location capacity and social network sizes (Alessandretti et al., 2018), incorporating a nested gravity model into the EPR model (Pappalardo et al., 2015), linking mobility to social ties and studying the dynamics of spatial choices based on social behavior (Toole et al., 2015), and integrating the circadian propensity of human mobility and Markov-based models into the EPR model (Jiang et al., 2016). In addition, the EPR model is also shown to be universal in human behavior modeling in general, such as user activities in recommendation systems (Onuma et al., 2009) and human behaviors in cyberspace (Hu et al., 2019). However, these models, often reliant on heuristic statistical assumptions, have limitations in accurately modeling the complex mobility patterns observed in real-world trajectories.
@@ 2.2. Synthetic Mobility Trajectory Benchmarks
The benchmarks for synthetic mobility trajectories can be classified into two categories (Kapp et al., 2023). First, statistical and distributional similarities are the most widely used benchmarks, such as Kullback-Leibler divergence (KLD) (Bindschaedler and Shokri, 2016), Jensen-Shannon divergence (JSD) (Ouyang et al., 2018; Feng et al., 2020), earth mover’s distance (EMD) (Anda et al., 2021), Root Mean Squared Error (RMSE) (Pappalardo and Simini, 2018), which are used to measure the similarities between real and generated trajectories in different aspects. For example, typical mobility statistics include the radius of gyration (Feng et al., 2020; Jiang et al., 2023; Yuan et al., 2022), the number of distinct locations visited per user per day (Feng et al., 2020; Xu et al., 2021), I-Rank (frequency of visiting personal top locations) (Feng et al., 2020; Pappalardo and Simini, 2018), and the number of daily trips per user (Pappalardo and Simini, 2018), trip lengths between consecutive trajectory points or between the origin and destination (Feng et al., 2020; Jiang et al., 2023; Xiong et al., 2023); spatial distributions characterize the distribution of locations based on factors like visits per location (i.e. G-Rank) (Feng et al., 2020; Jiang et al., 2023; Ouyang et al., 2018; Wang et al., 2021; Xiong et al., 2023) or location popularity ranking (Du et al., 2023); temporal distributions characterize the number of trips per hour of the day (Pappalardo and Simini, 2018), stay duration (Feng et al., 2020; Long et al., 2023; Ouyang et al., 2018) and time intervals between check-ins (Yuan et al., 2022). However, while these similarity metrics provide insights into the differences between real and generated data from various perspectives, they cannot fully reflect the ultimate utility of generated trajectories in supporting downstream tasks.
Second, benchmarking on downstream tasks recently emerged as an evaluation scheme for synthetic mobility trajectories. These tasks include road map updating (Cao and Li, 2021), next location prediction (Jiang et al., 2023; Long et al., 2023), and spreading simulation (Xiong et al., 2023; Feng et al., 2020; Yuan et al., 2022). However, these works often use a heuristically designed downstream task with one specific technique/algorithm to solve the task, which leads to unknown biases in the utility evaluation. As evidenced in our experiments in Appendix A, the performance of different techniques solving the same task varies; heuristically choosing the results of one technique as the benchmark is thus untrustworthy. Therefore, we propose a comprehensive task-based evaluation protocol to systematically benchmark synthetic mobility trajectories.
@@ 3.2.2. TPP Parameterization
Traditional TPPs often specify a simple parametric intensity function capturing relatively simple patterns in event occurrences, such as the Hawkes process (Hawkes, 1971), which often leads to poor results because of its limited flexibility in modeling complex data. In this context, neural TPPs are developed to use neural networks to learn complex dependencies from the history of TPPs, and then approximate the intensity function by some parametric forms (Shchur et al., 2021). However, these intensity-based approaches cannot achieve flexibility, efficiency, and ease-to-use simultaneously (Shchur et al., 2019). Alternatively, an intensity-free learning model of neural TPPs is proposed to use a mixture of log-normal distributions to represent the conditional probability density function $p^{*}(\tau)$ directly, and thus bypass the cumbersome intensity function (Shchur et al., 2019). Due to its advantage of closed-form sampling and likelihood computation, the intensity-free approach has been widely adopted in learning neural TPPs (Xu et al., 2021; Feng et al., 2023). We also adopt this approach in our work.
@@ 4.3.3. Location decoder with neural EPR
Location sampling in the return mode. In the return mode $z=return$, we consider the temporal periodicity of human behaviors, where individuals tend to revisit previously visited locations under a regular temporal distance, such as returning home daily (with a temporal distance of 24 hours) (Song et al., 2010a). Therefore, the return mode is designed to imitate such revisiting behaviors. Specifically, instead of imposing specific constraints or predefined temporal periodicity as (Yang et al., 2020a; Deng et al., 2023), we use temporal distance embeddings to learn the probability distribution of returning to previously visited locations. We define the temporal distance under an hour granularity between the $(i+1)$-th and $j$-th events as $\Delta^{t}_{i+1,j}=|t_{i+1}-t_{j}|\in\mathbb{Z}_{\geq 0}$. Subsequently, for the next event time $t_{i+1}$, we obtain its temporal distances to all previous events as follows:
@@
Note that here $p^{*}(l_{i+1}|\tau_{i+1},return)$ is not conditioned on the sampled next event category $k_{i+1}$ because the revisiting behavior depends only on the next event time. We then sample a previously visited location $l$ according to $p^{*}(l_{i+1}|\tau_{i+1},return)$. Our experiments also reveal that this learnt probability indeed exhibits a clear daily return pattern over different time distances (see Figure 4).
@@ 4.4.1. Model Training
In the training process, the above conditional probability distribution is conditioned on the real data without sampling. After training this objective function, we then train the user VAE on the learnt individual preference embeddings. The complexity of our model is discussed in Appendix B.
@@ 5.1.2. Baselines
We consider the following state-of-the-art baselines of three categories: statistical models Semi-Markov (Korolyuk et al., 1975) and TimeGeo (Jiang et al., 2016); neural TPP RMTPP (Du et al., 2016), ERTPP (Xiao et al., 2017), THP (Zuo et al., 2020), ActSTD (Yuan et al., 2022) and LogNormMix (Shchur et al., 2019); deep learning generative models LSTM (Huang et al., 2015), SeqGAN (Yu et al., 2017), MoveSim (Feng et al., 2020), VOLUNTEER (Long et al., 2023) and DiffTraj (Zhu et al., 2024). The baseline details are in Appendix C.
@@ 5.1.3. Statistical and Distributional Similarity Metrics
We adopt five popular metrics (Feng et al., 2020; Ouyang et al., 2018) to evaluate the resemblance between real and generated trajectories in different aspects. Distance measures the distance between successive locations in a trajectory. Radius of gyration is calculated as the root mean squared distance of all locations from the central one in a trajectory. Interval is computed as time intervals between successive events in a trajectory. DailyLoc computes the unique locations visited by users. Category computes the overall distribution of the POI categories. We use the Jensen-Shannon divergence (JSD) (Fuglede and Topsoe, 2004) as the similarity metric between the distributions of real and generated trajectories.